OpenStudy (anonymous):

Need help now with equations please: Solving for "n" 1) 5.7n + 4.1 =2.96 Solve for n 2) n8+ 1/2= -3 1/2 Thank u!! :)

OpenStudy (the_fizicx99):

5.7n + 4.1 =2.96 -4.1 -4.1 5.7n = -1.14 /5.7 /5.7 n = -0.2 Check, 5.7(-0.2) + 4.1 = 2.96 -1.14 + 4.1 = 2.96 2.96 = 2.96 n does equal -1.14

OpenStudy (the_fizicx99):

n8+ 1/2= -3 1/2 -1/2 -1/2 -3*2 = -6 + 1 = 5 5/2 - 1/2 4/2 = 2 n8 = 2 /8 /8 n = 0.25 ---------- 0.25(8) = 2 2 + 0.50 = 2.50 2.50 = 2.50 (-3 1/2 = 2.50)

OpenStudy (anonymous):

-1.14 is not correct, here are options for answers:

OpenStudy (the_fizicx99):

That was a typo, I meant to put 0.2

OpenStudy (anonymous):

-0.2, -3.2,-32, and -5

OpenStudy (the_fizicx99):

-0.2 * I copied and pasted from my work to not write it again. >.< its -0.2

OpenStudy (the_fizicx99):

You can check if you want by plugging in -0.2 for n

OpenStudy (anonymous):

will do

OpenStudy (anonymous):

thanks

OpenStudy (the_fizicx99):

Yw

OpenStudy (anonymous):

what about second one?

OpenStudy (anonymous):

Solution for 5.7n+4.1=2.96 equation: Simplifying 5.7n + 4.1 = 2.96 Reorder the terms: 4.1 + 5.7n = 2.96 Solving 4.1 + 5.7n = 2.96 Solving for variable 'n'. Move all terms containing n to the left, all other terms to the right. Add '-4.1' to each side of the equation. 4.1 + -4.1 + 5.7n = 2.96 + -4.1 Combine like terms: 4.1 + -4.1 = 0.0 0.0 + 5.7n = 2.96 + -4.1 5.7n = 2.96 + -4.1 Combine like terms: 2.96 + -4.1 = -1.14 5.7n = -1.14 Divide each side by '5.7'. n = -0.2 Simplifying n = -0.2

OpenStudy (anonymous):

wow, how did u do that ??

OpenStudy (the_fizicx99):

http://www.geteasysolution.com/5.7n+4.1=2.96 -sigh-

OpenStudy (the_fizicx99):

You can use wolfram alpha like i do or mathway for a better guide xD

OpenStudy (anonymous):

lol like i said, I do the problem in my paper and look up the same solution as mine and paste it here cause im too lazy to transfer it :P

OpenStudy (anonymous):

thank u all!!!

OpenStudy (the_fizicx99):

Yw